OpenStudy (bhavya):

find the speed of a stone which possesses kinetic energy of 20 joules and mass of 100 grams

OpenStudy (anonymous):

E = (1/2)mv², u know how to use this equation right?

OpenStudy (anonymous):

20 j = (1/2)(0.1kg)(v²) v=sqrt( 2(20 j) / (0.1kg) ) =srrt(400)m/s = 20m/s